For the following chemical equation, a student reacts 115 grams of pentane (C5H12)with 325 grams of oxygen. Calculate the mass of pentane and the mass of oxygen (in grams) that remained after the reaction was complete. Also,if the reaction produces a total of 2575 kJ of energy, calculate the percent yield.

C5H12(g) + 8 O2(g)5 CO2(g) + 6 H2O(g) + 3525kJ


1
Expert's answer
2021-03-26T05:45:57-0400

I) To know the mass of oxygen and pentane that remained after the reaction was completed, let's find the limiting reagent that was used up first.

From the balanced equation,

72g of pentane reacted with 256g of oxygen

115g of pentane should react with 256/72 x 115=409g of oxygen

But only 325g of oxygen was available, oxygen was therefore the limiting reagent and it was completely used up before the reaction was completed.

Now, let's find the mass of pentane that reacted with 325g of O2

256g of O2 reacted with 72g of pentane

325g of oxygen would react with 72/256 x 325 = 91g


Mass of pentane remaining= 115-91= 24g of pentane

Mass of oxygen remaining= 0g of oxygen


II) for the percentage yield of energy

%yield= experimental yield/theoretical yield x 100%

= 2575KJ/3525KJ x 100%=73%